Can someone help me solve this problem?
Rose's garden is in the shape of a trapezoid. If the height of the trapezoid is 16 m, one base is 20 m, and the area is 224 m^2, find the length of the other base.
Thanks,
Sher

Answer by Osran_Shri(18)


Area of a Trapezoid is given by:
Area = (1/2)*( base1 + Base2)* Height
Area = 224 m^2
Base1 = 20 m
Height = 16 m
Base2 = To be found out
224 = (1/2) * (20 + Base2)*16
20+Base2 = (224 * 2) / 16
20+Base2 = 448 / 16 = 28
Answer : Base2 = 28 - 20 = 8 m
Please contact me if you require any clarifications.
osran_con@yahoo.co.in